# Stochastic dominance for super heavy-tailed random variables

### v1

@@ Abstract
We introduce a class of super heavy-tailed distributions and establish the inequality that any weighted average of independent and identically distributed super heavy-tailed random variables stochastically dominates one such random variable. We show that many commonly used extremely heavy-tailed (i.e., infinite-mean) distributions, such as the Pareto, Fréchet, and Burr distributions, belong to the class of super heavy-tailed distributions. The established stochastic dominance relation is further generalized to allow negatively dependent or non-identically distributed random variables. In particular, the weighted average of non-identically distributed random variables stochastically dominates their distribution mixtures. Applications of these results in portfolio diversification, goods bundling, and inventory management are discussed. Remarkably, in the presence of super heavy-tailedness, the results that hold for finite-mean models in these applications are flipped.
@@ 1 Introduction
Distributions with infinite mean are ubiquitous in the realm of banking and insurance, and they are particularly useful in modeling catastrophic losses (Ibragimov et al. (2009)), operational losses (Moscadelli, (2004)), costs of cyber risk events (Eling and Wirfs, (2019)), and financial returns from technology innovations (Silverberg and Verspagen, (2007)); see also Chen et al. (2024b) for a list of empirical examples of distributions with infinite mean.
As the world is arguably finite (e.g., any loss is bounded by the total wealth in the world), why should we use models with infinite mean as mathematical tools? There are perhaps two reasons. The first reason is that, after careful statistical analysis, infinite-mean models often fit extremely heavy-tailed datasets better than finite-mean models. The second reason is that the sample mean of iid samples of heavy-tailed data may not converge or may even tend to infinity as the sample size increases. Therefore, it is not sufficient to conclude that infinite-mean models are unrealistic by the finiteness of the sample mean. Indeed, models with infinite moments are not “improper" as emphasized by Mandelbrot, (1997), and they have been extensively used in the financial and economic literature (see Mandelbrot, (1997) and Cont (2001)).
This paper focuses on establishing some stochastic dominance relations for infinite-mean models, applicable in several contexts. For two random variables $X$ and $Y$, $X$ is said to be stochastically smaller than (or stochastically dominated by) $Y$, denoted by $X\leq_{\rm st}Y$, if $\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)$ for all $x\in\mathbb{R}.$ The order $\leq_{\rm st}$ is arguably the strongest form of commonly used stochastic orders; see Levy (2016) for their applications in decision making and Müller and Stoyan, (2002) and Shaked and Shanthikumar (2007) for the mathematics. Let $X$ be a Pareto random variable with infinite mean and $X_{1},\dots,X_{n}$ be iid copies of $X$. For a nonnegative vector $\left(\theta_{1},\dots,\theta_{n}\right)$ with $\sum_{i=1}^{n}\theta_{i}=1$, Chen et al. (2024a) showed that
@@
The inequality is strict if at least two components of $\left(\theta_{1},\dots,\theta_{n}\right)$ are positive. The case of two Pareto random variables with tail parameter 1/2 has been studied in Example 7 of Embrechts et al., (2002); see Section 3 for the precise definition of the Pareto distribution. A closely related result of Ibragimov, (2005) implies that (1) also holds if $X_{1},\dots,X_{n}$ are iid positive one-sided stable random variables with infinite mean.
@@
Inequality (1) provides very strong implications in decision making as it surprisingly holds in the strongest form of risk comparison. Intuitively, if $X_{1},\dots,X_{n}$ are treated as losses in a portfolio selection problem, any agent who prefers less loss will choose to take one of $X_{1},\dots,X_{n}$ instead of allocating their risk exposure over different losses. Other applications of (1) include optimal bundling problems (Ibragimov and Walden, (2010)), risk management (Chen et al. (2024a)), and risk sharing (Chen et al. (2024b)). It is worth noting that (1) cannot hold if at least two components of $\left(\theta_{1},\dots,\theta_{n}\right)$ are positive and $X_{1},\dots,X_{n}$ have finite mean (see Proposition 2 of Chen et al. (2024a)).
@@
Theorem 1 of Chen et al. (2024a) shows that (1) also holds for weakly negatively associated super-Pareto random variables $X_{1},\dots,X_{n}$. The class of super-Pareto random variables is quite broad and can be obtained by applying increasing and convex transforms to a Pareto random variable with tail parameter $1$. Important examples of super-Pareto distributions include the Pareto, generalized Pareto, Burr, paralogistic, and log-logistic distributions, all with infinite mean.
Given the applicability of (1), this paper aims to further generalize the inequality in two aspects: the marginal distribution and the dependence structure of $(X_{1},\dots,X_{n})$. We will first show that (1) holds for a rather large class of distributions, which we call super heavy-tailed (Theorem 1). The class of super heavy-tailed distributions includes super-Pareto distributions and has several nice properties (Propositions 2 and 3), leading to a very rich class of models; Section 3 gives many examples of super heavy-tailed distributions.
With a slight modification of the proof of Theorem 1, we show in Theorem 2 that (1) holds for negatively lower orthant dependent (Block et al., (1982)) super heavy-tailed random variables $X_{1},\dots,X_{n}$. Negative lower orthant dependence has been widely used in statistics and operations research and it relates to several popular notions of negative dependence in the literature such as negative association (Alam and Saxena (1981) and Joag-Dev and Proschan (1983)). As the class of super heavy-tailed distributions includes super-Pareto distributions and negative lower orthant dependence is implied by weak negative association, Theorem 2 implies Theorem 1 (i) of Chen et al. (2024a). Remarkably, while Theorem 2 is more general, it is shown by a much more concise proof than Chen et al. (2024a).
@@
We proceed to study (1) given non-identically distributed random variables $X_{1},\dots,X_{n}$. Since $X_{1},\dots,X_{n}$ do not follow the same distribution, the choice of $X$ becomes unclear. A possible choice is to let $X$ follow the generalized mean of the distributions of $X_{1},\dots,X_{n}$; note that the generalized mean of distribution functions is a distribution function. A special case is the arithmetic mean, which leads to the commonly used distribution mixture models. Considering a sufficiently large class of distributions, Theorem 3 shows that (1) holds if the distribution of $X$ is the generalized mean with non-negative power of the distributions of $X_{1},\dots,X_{n}$. To our best knowledge, Theorem 3 is the first attempt to establish a non-trivial version of (1) for non-identically distributed random variables.
The rest of the paper is organized as follows. In Section 2, we present some first observations on (1). We introduce the class of super heavy-tailed distributions and establish our main result (Theorem 1) in Section 3. Section 4 presents several generalizations of Theorem 1 to negative dependence and non-identical marginal distributions. Several applications of (1) are briefly discussed in Section 5. Section 6 concludes the paper. The appendix contains the proof of Proposition 5.
@@ 1.1 Notation, conventions and definitions
A function $f$ on $(0,\infty)$ is said to be sub-additive if $f(x+y)\leq f(x)+f(y)$ for any $x,y>0$. If the inequality is strict, we say $f$ is strictly sub-additive. For a random variable $X\sim F$, denote by $\operatorname*{ess\text{-}inf}X$ ($\operatorname*{ess\text{-}inf}F$) and $\operatorname*{ess\text{-}sup}X$ ($\operatorname*{ess\text{-}sup}F$) its essential infimum and essential supremum. Denote by $\Delta_{n}$ the standard simplex, that is, $\Delta_{n}=\{\bar{\theta}\in[0,1]^{n}:\sum_{i=1}^{n}\theta_{i}=1\}$, where we use notation $\bar{\theta}$ for a vector $(\theta_{1},\dots,\theta_{n})$. We will also use $[n]$ to denote the set of indices $1,\dots,n$.
@@
We say that a random variable $X$ is stochastically smaller than (or stochastically dominated by) a random variable $Y$, if
$\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)\quad\text{for all }x\in\mathbb{R}.$
We use notation $X\leq_{\rm st}Y$ for this property and refer to this order as stochastic order; see Shaked and Shanthikumar (2007) for properties of stochastic order. We write $X<_{\rm st}Y$ if $\mathbb{P}(X\leq x)>\mathbb{P}(Y\leq x)$ for all $x>\operatorname*{ess\text{-}inf}X$.
@@ 2 Some observations on the stochastic dominance
$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n},$ (2)
for all $\bar{\theta}\in\Delta_{n}$, where $X_{1},\dots,X_{n}$ are iid copies of $X$. We will also say that a distribution $F$ satisfies property (2) if a random variable $X\sim F$ satisfies it.
Since (2) holds if a constant is added to $X$, we will, without loss of generality, only consider random variables with essential infimum 00. To avoid trivialities, we also assume throughout that $n\geq 2$ and $\theta_{i}>0$ for all $i\in[n]$. We will also be interested in distributions, and random variables, for which property (2) holds with a strict inequality. Let us start by formulating and providing some straightforward observations of (2).
@@
Assume that random variables $X$ and $Y$ satisfy property (2) and are independent. Then the following statements hold.
@@
    A random variable $aX+b$ with $a\geq 0$ and $b\in\mathbb{R}$ satisfies (2).
@@
    Random variables $\max\{X,c\}$ and $\max\{X,Y\}$ satisfy (2), with $c\geq 0$.
@@
    A random variable $g(X)$ with a convex non-decreasing function $g$ satisfies (2). In addition, if $X$ satisfies (2) with a strict inequality, $g$ is convex and strictly increasing, then $g(X)$ also satisfies (2) with a strict inequality.
@@
    Proposition 2 of Chen et al. (2024a) shows that if $X,X_{1},\dots,X_{n}$ are identically distributed with finite mean, (2) holds if and only if $X_{1}=\dots=X_{n}$ almost surely. The desired result follows.
@@
    We will prove only the stronger property for the maximum of two random variables. Let $X_{1},\dots,X_{n}$ be iid copies of $X$, $Y_{1},\dots,Y_{n}$ be iid copies of $Y$, and $\{X_{i}\}_{i\in[n]}$ and $\{Y_{i}\}_{i\in[n]}$ be independent. For $x\in\mathbb{R}$ and $\bar{\theta}\in\Delta_{n}$, we have
@@
    Since $g$ is convex and non-decreasing, $g(X)\leq_{\rm st}g(\sum_{i=1}^{n}\theta_{i}X_{i})\leq\sum_{i=1}^{n}\theta_{i}g(X_{i})$, where the first inequality holds as stochastic order is preserved under monotone transforms and the second inequality is to be understood in the almost sure (and therefore also stochastic) sense and is due to convexity of $g$. ∎
Remark 1.
Properties (ii)-(iv) above demonstrate that, even if one knows only several random variables satisfying (2), it is possible to construct many more. Of special interest is property (iii), which does not require any specific distributional properties of $X$ and $Y$ apart from property (2).
3 Super heavy-tailed distributions and stochastic dominance
In this section, we introduce a class of distributions we refer to as super heavy-tailed. We then prove that all such distributions satisfy property (2). We also explore the properties of these distributions and demonstrate that many well-known distributions with infinite mean belong to this class. Along with the results of Proposition 1, this shows that the class of distributions satisfying property (2) is large.
@@ 3 Super heavy-tailed distributions and stochastic dominance
Let $F$ be a distribution function with $\operatorname*{ess\text{-}inf}F=0$ and $h_{F}(x)=-\log F\left(1/x\right)$ for $x\in(0,\infty)$. If $h_{F}$ is (strictly) sub-additive, we say $F$ is (strictly) super heavy-tailed. We also say a random variable $X\sim F$ is (strictly) super heavy-tailed if $F$ is (strictly) super heavy-tailed.

### v3

# Risk aggregation and stochastic dominance for a class of heavy-tailed distributions

@@ Abstract
We introduce a new class of heavy-tailed distributions for which any weighted average of independent and identically distributed random variables is larger than one such random variable in (usual) stochastic order. We show that many commonly used extremely heavy-tailed (i.e., infinite-mean) distributions, such as the Pareto, Fréchet, and Burr distributions, belong to this class. The established stochastic dominance relation can be further generalized to allow negatively dependent or non-identically distributed random variables. In particular, the weighted average of non-identically distributed random variables dominates their distribution mixtures in stochastic order.
@@ 1 Introduction
Distributions with infinite mean are ubiquitous in the realm of banking and insurance, and they are particularly useful in modeling catastrophic losses (Ibragimov et al. (2009)), operational losses (Moscadelli, (2004)), costs of cyber risk events (Eling and Wirfs, (2019)), and financial returns from technology innovations (Silverberg and Verspagen, (2007)); see also Chen and Wang (2025) for a list of empirical examples of distributions with infinite mean.
As the world is arguably finite (e.g., any loss is bounded by the total wealth in the world), why should we use models with infinite mean as mathematical tools? The main reason is that infinite-mean models often fit extremely heavy-tailed datasets better than finite-mean models. Moreover, the sample mean of iid samples of heavy-tailed data may not converge or may even tend to infinity as the sample size increases. Therefore, it is not sufficient to conclude that infinite-mean models are unrealistic by the finiteness of the sample mean. Indeed, models with infinite moments are not “improper" as emphasized by Mandelbrot, (1997), and they have been extensively used in the financial and economic literature (see Mandelbrot, (1997) and Cont (2001)).
This paper focuses on establishing some stochastic dominance relations for infinite-mean models. For two random variables $X$ and $Y$, $X$ is said to be smaller than $Y$ in stochastic order, denoted by $X\leq_{\rm st}Y$, if $\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)$ for all $x\in\mathbb{R}$; see Müller and Stoyan, (2002) and Shaked and Shanthikumar (2007) for extensive accounts of properties of stochastic dominance. Let $X$ be a positive one-sided stable random variable with infinite mean and $X_{1},\dots,X_{n}$ be iid copies of $X$. For a nonnegative vector $\left(\theta_{1},\dots,\theta_{n}\right)$ with $\sum_{i=1}^{n}\theta_{i}=1$, Ibragimov, (2005) showed that
@@
Recently, Arab et al., (2024), Chen et al. (2024a), and Müller, (2024) have shown that inequality (1) holds for more general classes of distributions. The case of two Pareto random variables with tail parameter 1/2 was studied in Example 7 of Embrechts et al., (2002); see Section 3 for the precise definition of the Pareto distribution.
@@
Inequality (1) provides very strong implications in decision making as it surprisingly holds in the strongest form of risk comparison. If $X_{1},\dots,X_{n}$ are treated as losses in a portfolio selection problem, any agent who prefers less loss will choose to take one of $X_{1},\dots,X_{n}$ instead of allocating their risk exposure over different losses. This observation is counterintuitive, contrasting with the common belief that diversification reduces risk. Other applications of (1) include optimal bundling problems (Ibragimov and Walden, (2010)) and risk sharing (Chen et al. (2024b)).
@@
In this paper, we will study (1) where $X_{1},\dots,X_{n}$ are possibly negatively dependent, a case not considered in Ibragimov, (2005), Arab et al., (2024), and Müller, (2024). Chen et al. (2024a) have shown that (1) also holds for weakly negatively associated super-Pareto random variables $X_{1},\dots,X_{n}$. The class of super-Pareto random variables is quite broad and can be obtained by applying increasing and convex transforms to a Pareto random variable with tail parameter $1$. Examples of super-Pareto distributions include the Pareto, generalized Pareto, Burr, paralogistic, and log-logistic distributions, all with infinite mean.
This paper aims to further generalize the result of Chen et al. (2024a) in two aspects: the marginal distribution and the dependence structure of $(X_{1},\dots,X_{n})$. In Section 3, we first introduce a new class of distributions, which has several nice properties (Propositions 2 and 3) and includes the class of super-Pareto distributions as a special case. Within this class of distributions, we show in Theorem 1 that (1) holds for identically distributed random variables $X_{1},\dots,X_{n}$ that are negatively lower orthant dependent (Block et al., (1982)). It is well known that the behavior of the sum of extremely heavy-tailed random variables is dominated by the maximum of the summands (see Embrechts et al. (1997)). Therefore, a possible reason why (1) is preserved when transitioning from independence to negative dependence is because under negative dependence, random variables that take small to moderate values will push the other random variables to take large values with a larger probability, leading to a stochastically larger $\sum_{i=1}^{n}\theta_{i}X_{i}$. The situation is different for positively dependent random variables; see Remark 5. As negative lower orthant dependence is more general than weak negative association, Theorem 1 (i) of Chen et al. (2024a) is implied by Theorem 1. Remarkably, while Theorem 1 is more general, it is shown by a much more concise proof.
@@
In Section 4, we proceed to study (1) given non-identically distributed random variables $X_{1},\dots,X_{n}$. Since $X_{1},\dots,X_{n}$ do not follow the same distribution, the choice of $X$ becomes unclear. A possible choice is to let $X$ follow the generalized mean of the distributions of $X_{1},\dots,X_{n}$. A special case is the arithmetic mean, which leads to the commonly used distribution mixture models. Considering a rather large class of distributions, Theorem 2 shows that (1) holds if the distribution of $X$ is the generalized mean with non-negative power of the distributions of $X_{1},\dots,X_{n}$. To our best knowledge, Theorem 2 is the first attempt to establish a non-trivial version of (1) for non-identically distributed random variables.
The rest of the paper is organized as follows. In Section 2, we present some first observations on (1). Sections 3 and 4 present the main results. Section 5 compares our results with the literature. Section 6 concludes the paper. The appendix contains the proofs of Propositions 2 and 3 as well as some examples in the new class of distributions.
@@ 1.1 Notation, conventions and definitions
A function $f$ on $(0,\infty)$ is said to be subadditive if $f(x+y)\leq f(x)+f(y)$ for any $x,y>0$. If the inequality is strict, we say $f$ is strictly subadditive. For a random variable $X\sim F$, denote by $\operatorname*{ess\text{-}inf}X$ ($\operatorname*{ess\text{-}inf}F$) and $\operatorname*{ess\text{-}sup}X$ ($\operatorname*{ess\text{-}sup}F$) its essential infimum and essential supremum. Denote by $\Delta_{n}$ the standard simplex, that is, $\Delta_{n}=\{\bar{\theta}\in[0,1]^{n}:\sum_{i=1}^{n}\theta_{i}=1\}$, where we use notation $\bar{\theta}$ for a vector $(\theta_{1},\dots,\theta_{n})$. Let $\Delta_{n}^{+}=\Delta_{n}\cap(0,1)^{n}$. We will also use $[n]$ to denote the set of indices $1,\dots,n$. For a distribution function $F$, its generalized inverse is defined as
$F^{-1}(p)=\inf\{t\in\mathbb{R}:F(t)\geq p\},~{}p\in(0,1).$
@@
We say that a random variable $X$ is smaller than a random variable $Y$ in stochastic order, denoted by $X\leq_{\rm st}Y$, if $\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)$ for all $x\in\mathbb{R}$. For random variables $X$ and $Y$ with support $[c,\infty)$ where $c\in\mathbb{R}$, we write $X<_{\rm st}Y$ if $\mathbb{P}(X\leq x)>\mathbb{P}(Y\leq x)$ for all $x>c$.
@@ 2 Some observations on the stochastic dominance
$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}\mbox{~{}for all $\bar{\theta}\in\Delta_{n}$,}$ (SD)
where $X_{1},\dots,X_{n}$ are independent or negatively dependent with the marginal laws equal to $X$ (see Section 3.2 for the precise definition of negative dependence). We will also say that a distribution $F$ satisfies property (SD) if a random variable $X\sim F$ satisfies it. If some of $\theta_{1},\dots,\theta_{n}$ are 0, we can simply reduce the dimension of our problem. Therefore, for most of our results, we will assume $\bar{\theta}\in\Delta_{n}^{+}$.
Since (SD) holds if a constant is added to $X$, we will, without loss of generality, only consider random variables with essential infimum 00. We will also be interested in distributions, and random variables, for which property (SD) holds with a strict inequality. Let us start by formulating and providing some straightforward observations of (SD).
@@
Assume that random variables $X$ and $Y$ satisfy property (SD) and are independent. Then the following statements hold.
@@
    A random variable $aX+b$ with $a\geq 0$ and $b\in\mathbb{R}$ satisfies (SD).
@@
    Random variables $\max\{X,c\}$ and $\max\{X,Y\}$ satisfy (SD), with $c\geq 0$.
@@
    A random variable $g(X)$ with a convex non-decreasing function $g$ satisfies (SD). In addition, if $X$ satisfies (SD) with a strict inequality, $g$ is convex and strictly increasing, then $g(X)$ also satisfies (SD) with a strict inequality.
@@
    This is implied by Proposition 2 of Chen et al. (2024a).
@@
    We will prove only the stronger property for the maximum of two random variables. Let $X_{1},\dots,X_{n}$ follow the distribution of $X$, $Y_{1},\dots,Y_{n}$ follow the distribution of $Y$, and $\{X_{i}\}_{i\in[n]}$ and $\{Y_{i}\}_{i\in[n]}$ be independent. For $x\in\mathbb{R}$ and $\bar{\theta}\in\Delta_{n}$, we have
@@
    Since $g$ is convex and non-decreasing, $g(X)\leq_{\rm st}g(\sum_{i=1}^{n}\theta_{i}X_{i})\leq\sum_{i=1}^{n}\theta_{i}g(X_{i})$, where the first inequality holds as stochastic order is preserved under non-decreasing transforms and the second inequality is to be understood in the almost sure (and therefore also stochastic) sense and is due to convexity of $g$. ∎
Properties (ii)-(iv) above demonstrate that, even if one knows only several random variables satisfying (SD), it is possible to construct many more. Of special interest is property (iii), which does not require any specific distributional properties of $X$ and $Y$ apart from property (SD).
3 A class of heavy-tailed distributions and stochastic dominance
In this section, we introduce a new class of heavy-tailed distributions. We explore several properties of this class and demonstrate that it contains many well-known distributions with infinite mean. We then prove that all distributions in this class satisfy property (SD). Along with the results of Proposition 1, this shows that the class of distributions satisfying property (SD) is large.
3.1 A class of heavy-tailed distributions
@@
Let $F$ be a distribution function with $\operatorname*{ess\text{-}inf}F=0$ and let $h_{F}(x)=-\log F(1/x)$ for $x\in(0,\infty)$. We say that $F$ belongs to $\mathcal{H}$, denoted by $F\in\mathcal{H}$, if $h_{F}$ is subadditive. We write $F\in\mathcal{H}_{s}$ if $h_{F}$ is strictly subadditive. For $X\sim F$, we also write $X\sim\mathcal{H}$ (resp. $X\sim\mathcal{H}_{s}$) if $F\in\mathcal{H}$ (resp. $F\sim\mathcal{H}_{s}$).
Remark 1.
By properties of subadditive functions (e.g., Theorem 7.2.4 and Theorem 7.2.5 of Hille and Phillips, (1996)), $F\in\mathcal{H}$ if $h_{F}(x)/x$ is decreasing or $h_{F}$ is concave.
In the case of continuous distribution $F$, $F\in\mathcal{H}$ holds if and only if the survival function of $1/X$ is log-superadditive where $X\sim F$. We will see later that all distributions in $\mathcal{H}$ have infinite mean and because of that we say $\mathcal{H}$ is a class of heavy-tailed distributions. Note that the definition of heavy-tailed distributions varies in different contexts; see, e.g., Remark 3. Below are some examples in class $\mathcal{H}$.
Example 1 (Fréchet distribution).
For $\alpha>0$, the Fréchet distribution, denoted by Fréchet$(\alpha)$, is defined as
$F(x)=\exp(-x^{-\alpha}),~{}~{}x>0.$
If $\alpha\leq 1$, $F$ has infinite mean. It is easy to check that $F\in\mathcal{H}$ if $\alpha\leq 1$ and $F\in\mathcal{H}_{s}$ if $\alpha<1$, since for any $x,y>0$,
As $h_{F}$ is additive when $\alpha=1$, Fréchet$(1)$ distribution can be thought as a “boundary” of class $\mathcal{H}$.
Example 2 (Pareto$(1)$ distribution).
$F(x)=1-\frac{1}{(x+1)^{\alpha}},~{}~{}x>0.$
Pareto$(\alpha)$ distributions have infinite mean if $\alpha\leq 1$. Taking second derivative of $h_{F}$ when $\alpha=1$, we have $h^{\prime\prime}_{F}(x)=-1/(x+1)^{2}.$ Hence $h_{F}$ is concave and ${\rm Pareto}(1)\in\mathcal{H}_{s}$.
We can show that Pareto$(\alpha)$ with $\alpha\leq 1$, as well as many other infinite-mean distributions in Table 1 also belong to $\mathcal{H}$ either directly using the definition, or using some closure properties of $\mathcal{H}$ in Propositions 2 and 3 provided below; see Appendix for detailed derivations of examples in Table 1 and the proofs of Propositions 2 and 3.
Distribution functions Parameters
Fréchet distribution $F(x)=\exp(-x^{-\alpha}),~{}~{}x>0$ $\alpha\leq 1$
Pareto distribution $F(x)=1-(x+1)^{-\alpha},~{}~{}x>0$ $\alpha\leq 1$
Generalized Pareto distribution $F(x)=1-\left(1+\xi(x/\beta)\right)^{-1/\xi},~{}~{}x>0$ $\xi\geq 1$
Burr distribution $F(x)=1-\left(x^{\tau}+1\right)^{-\alpha},~{}~{}x>0$ $\alpha,\tau\leq 1$
Inverse Burr distribution $F(x)=\left(x^{\tau}/(x^{\tau}+1)\right)^{\alpha},~{}~{}x>0$ $\alpha>0$, $\tau\leq 1$
Log-Pareto distribution $F(x)=1-(\log(x+1)+1)^{-\alpha},~{}~{}x>0.$ $\alpha\leq 1$
Stoppa distribution $F(x)=\left(1-(x+1)^{-\alpha}\right)^{\beta},~{}~{}x>0$ $\alpha\leq 1$, $\beta>0$
Table 1: Examples of distributions in $\mathcal{H}$
Proposition 2.
Let $X\sim F$ where $F\in\mathcal{H}$. The following statements hold.
  1. (i)
    If $F$ is strictly increasing on $[0,\infty)$ and $\mathbb{P}(X<\infty)=1$, then $F$ is continuous on $[0,\infty)$.
  2. (ii)
    For $\beta>0$, $F^{\beta}\in\mathcal{H}$.
  3. (iii)
    If, in addition, a random variable $Y\sim G$, where $G\in\mathcal{H}$, is independent of $X$, then $\max\{X,Y\}\in\mathcal{H}$. In terms of distribution functions, if $F,G\in\mathcal{H}$, then $FG\in\mathcal{H}$.
  4. (iv)
    For a non-decreasing, convex, and non-constant function $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ with $f(0)=0$, $f(X)\in\mathcal{H}$.
Proposition 3.
Let $\bar{\theta}\in\Delta_{n}^{+}$. If distribution functions $F_{1},\dots,F_{n}\in\mathcal{H}$ and $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{n}$, then $\sum_{i=1}^{n}\theta_{i}F_{i}\in\mathcal{H}$.
It is clear that the various transforms of distributions in Proposition 2 from our class generate many different distributions, showing that the class $\mathcal{H}$ is indeed rather large. Suppose that $F_{1},\dots,F_{n}$ are Pareto distributions with possibly different tail parameters $0<\alpha_{1},\dots,\alpha_{n}\leq 1$. As $F_{1},\dots,F_{n}$ are comparable in stochastic order, by Proposition 3, mixtures of $F_{1},\dots,F_{n}$ are in $\mathcal{H}$.
3.2 Negative lower orthant dependence
Definition 3 (Block et al., (1982)).
Negative lower orthant dependence includes independence as a special case. It is commonly used in various research areas and it is implied by other popular notions of negative dependence in the literature, such as negative association (Alam and Saxena (1981) and Joag-Dev and Proschan (1983)), negative orthant dependence (Block et al., (1982)), and negative regression dependence (Lehmann (1966) and Block et al., (1985)) see, e.g., Chi et al., (2024) for the implications of these notions.
3.3 Main result